Find two angles such that:
The angles are supplementary and the measure of the smaller is one-half the measure of the larger.

Found 2 solutions by Cromlix, macston:
Answer by Cromlix(4381)

Hi there,
Angles that are supplementary
add up to give 180 degrees.
120 degrees and 60 degrees?
Hope this helps:-)

Answer by macston(5194)

.
S=smaller angle; Supplement=180-S
S=1/2(180-S)
2S=180-S
3S=180
S=60 degrees
ANSWER 1: The measure of the smaller angle is 60 degrees.
Supplement=180-S=180-60=120 degrees
ANSWER 2: The supplement measures 120 degrees.